Formulate a system of equations for the situation below and solve A manufacturer of womens blouses makes three types of blouses: sleeveless, short-sleeve, and long-sleeve. The time (in minutes) required by each department to produce a dozen blouses of each type is shown in the following table Short Long- Sleeveless Sleeve Sleeve Cutting Sewing Packaging 12 24 8 15 28 8 The cutting, sewing, and packaging departments have available a maximum of 78, 160, and 48 labor-hours, respectively, per day. How many dozens of each type of blouse can be produced each day if the plant is operated at full capacity? sleeveless short-sleeve long-sleeve dozen dozen dozen
Answer #1

Tong Short sleeve 12 sleeve is. 28 22 Solved; Sleerelers Cutting Sewing 24 8 Packaging The cutting, sewing and packaging depa(D Nowi Minutee spent cutting sleevless shirt=9x Minutes spont cutting short sleeve shirts = 12y Minutes spent cutting longSo the speten of equcation we must solve is 922 +12y +153 = 4688 22x +2y +283 = 9600 +83 By Gauss Jordan Gliminations, 6x + s513 43 - 16/3 520 - 1840 263 -240 R2 - R2X-3 16 473 53 13/8 520 345 o -2 -240 RRI y R2 60 o 1318 345 o 6 -240 R3 - R3 = (-2)Ra R2 - 13 R3 - B 120 0 150 120 Therefore; x = 120 y = = 150 3= to x= Sleeveless shirts 120 y = - Short sleeve shirts = 150 L